# A model of thermophoresis of colloidal proteins in water using non-Fickian diffusion currents

### v1

@@ Abstract
In 1928, Chapman generalised Einstein’s theory of diffusion for non-uniform fluids to show the presence of a non-Fickian diffusion current, which he considered important in thermodiffusion (Ludwig-Soret effect). In 1941, Kiyosi Itô proposed the formal methods of stochastic calculus in the presence of spatially dependent diffusion, yielding the same non-Fickian diffusion current as shown by Chapman. The phenomenon of thermodiffusion and thermophoresis happens in the presence of a temperature gradient, which makes diffusion space-dependent. The role of solvation forces in thermophoresis will only be clearer once that of diffusion is understood properly. In this paper, we investigate the importance of Chapman’s non-Fickian diffusion current on the thermophoretic motion of colloidal particles in water (with weak salt concentration). We show that all the general features of variations of the Soret coefficient $S_{T}$ with temperature can be captured using Chapman’s non-Fickian diffusion current. We compare our theoretical results with experimental plots of the Soret coefficients for three polypeptides in aqueous solution: Lysozyme, BLGA, and Poly-L-Lysine, and find a strong match. We emphasise that, in addition to the yet-to-be-understood details of solvation forces, Chapman’s non-Fickian diffusion current is an indispensable element that needs to be taken into account for a complete understanding of thermophoresis and thermodiffusion.
Suggested keywords
@@ I INTRODUCTION
The thermophoretic motion of colloidal particles is closely related to thermodiffusion, first reported by Ludwig in 1859 [20] and subsequently observed by Soret in 1879 [21]. The process of thermodiffusion, also known as the Ludwig-Soret effect, involves partial segregation of the components of a solution or a fluid mixture in the presence of a thermal gradient. Thermal diffusion is understood to play a significant role in the convective separation of components in a fluid mixture. It is the interplay of convective flow and diffusion that underlie the thermodiffusive partial separation of the mixture’s components. Therefore, a proper theoretical understanding of these complex processes of thermodiffusion and thermophoresis could only have meaningfully begun not before the advent of the physical theory of diffusion based on Einstein’s work on Brownian motion in 1905 [22, 23]. As is well acknowledged, Einstein’s theory of Brownian motion provides the microscopic origin of Fickian diffusion. What lacks wider acknowledgement is that Einstein’s theory of Brownian motion also underlies the non-Fickian diffusion in inhomogeneous media and fields.
@@
The Fick’s law of diffusion in a homogeneous medium gives the Fickian diffusion current density (e.g., in one dimension along the $x$-direction), written as $j_{F}=-D(dC(x)/dx)$ where the diffusion constant is $D$. Diffusivity is a constant that indicates the space in which diffusion takes place is uniform. The non-Fickian diffusion current density $j_{NF}$ will arise in an inhomogeneous medium where the diffusivity $D(x)$ is a function of space. This current will arise as an additive part of the total diffusion current density j=dD(x)C(x)/dxj=-dD(x)C(x)/dx, where $j_{NF}=-C(x)(dD(x)/dx)$. At the origin of both the parts of the diffusion current $j_{F}$ and $j_{NF}$, the assumption of the local homogeneity and isotropy of space, which is a fundamental requirement of Einstein’s theory of Brownian motion, remains intact [28]. This is why these are diffusion currents, unlike drift currents that break the local isotropy of space. From various perspectives, many groups have studied enhanced diffusion in active and other systems like enzymes and also in equilibrium Brownian systems, where this Chapman-Itô non-Fickian diffusion current has been taken into account  [29, 30, 31, 32, 33, 28, 34, 35, 36]. Works by Di Pu et al. have considered the Chapman-Itô non-Fickian contribution in the field of thermophoresis  [37, 38, 39, 40]. Seuwin et al. [41] have tried to explain the sign change in protein thermophoresis from a hydrogen-bonding perspective, to cite a few.
@@
Keeping this general picture in mind, the process of thermophoresis (and thermodiffusion) can be looked at as a phenomenon arising from the competition of three current densities. (1) The Fickian diffusion current, (2) the Non-Fickian diffusion current, and (3) the drift current due to solvation forces. In the present paper, we demonstrate the interplay of these three competing ingredients, resulting in Soret coefficients observed in three distinct experiments on aqueous solutions of Lysozyme, BLGA, and Poly-L-Lysine. We chose these three experimental results to avoid cases involving strong electrostatic forces. We show that non-Fickian diffusion is a major contributor to thermophoresis and thermodiffusion. For the class of systems under consideration, we show that the local density of water (assumed to have low salt concentration), its thermal expansion coefficient, and a generic local pressure gradient can account for the observed $S_{T}(T)$ of all the test cases. We also show that our model can match the experimentally observed equilibrium saturation concentration curve for lysozyme quite well, underscoring the important role the non-Fickian current plays in equilibrium and nonequilibrium phenomena. Moreover, we show that the theoretical model built with these specific processes of thermophoresis in colloids in mind can also account for the general features of the Soret coefficient for thermodiffusion. Therefore, the model framework developed in this paper, based on results from kinetic theory and empirical inputs, may capture some fundamental aspects of transport driven by thermal gradients in a fluid.
@@ II.1 The general structure of the diffusivity
The $D_{0}(T)$ represents the bare single-particle diffusivity, capturing the explicit temperature dependence arising from solute–solvent interactions (e.g. viscosity, hydration, and thermal agitation) in the dilute limit. The concentration-dependent factor of diffusivity $\Lambda(C)$ accounts for the dependence of diffusivity on crowding and ensuing hydrodynamic effects in general [42]. The parameter $b$ sets the strength of the concentration dependence of the diffusivity. One may imagine that there exists a scale of concentration $C_{0}$, where $b=C_{0}^{-\beta}$. The dependence of the diffusivity on the local concentration is dominant for $C>C_{0}$.
@@
At this stage, we can define two dominant regimes of the concentration dependence of the diffusivity as (1) the weak concentration $bC^{\beta}<<1$ and (2) the strong concentration $bC^{\beta}>>1$, where the corresponding Soret coefficients at the leading order are, respectively,
@@ II.2 Temperature dependence of the dynamic viscosity
Thermal gradients can alter the local solvent density, thereby modifying the local dynamic viscosity through a range of processes, e.g., from local hydrodynamics to hydrogen-bonding chemistry. The strength of this viscosity modification due to temperature can be quantified by the solvent thermal expansion coefficient $\alpha_{w}(T)$, defined in terms of the temperature-dependent solvent density $\rho_{w}(T)$ as
@@ III Results: Comparison with experiment
Let us compare our model with the curves fitted to the experimental data for hen egg-white lysozyme and other polypeptide solutions reported in ref. [43, 49, 50]. The empirical fitting function chosen to fit the experimentally obtained Soret coefficient is
@@
where $S^{\infty}_{T}$ is the high-temperature thermophobic limit, $T^{*}$ is the temperature at which $S^{\prime}_{T}$ switches sign, and $T_{0}$ denotes the scale of exponential growth. Note that the transition temperature $T^{*}$ is poorly sensitive to salt addition and Lysozyme charge [43, 49, 50]. The above-mentioned empirical function for the Soret coefficient in these systems is applicable over a wide range of pH and ionic strength, indicating that neither pH nor ionic strength significantly affects the ansatz. Our model at present does not include explicit contributions from pH or ionic strength; beyond the limit which can be accounted for through the parameters $c$ and $a$. Thus, our model is well-suited for comparison with experimental data that, in general, matches the empirical Soret coefficient of eq.(16). The value of the three free parameters of $S_{T}^{\prime}$ in eq.(16) and the corresponding three free parameters in the expression of $S_{T}$ in our model, which make the curves match $S^{\prime}_{T}$ that matches the experimental data, are shown in Table 1. Figure 2 compares curves with the $S_{T}$ from our model plotted as dots. The experiments on lysozyme, whose data are presented here, were carried out in a sodium acetate buffer at pH=4.65=4.65, with a lysozyme concentration $7~\mathrm{g\,L^{-1}}$ and $0.4~\mathrm{M}$ NaCl (corresponding to $2.3\%~\mathrm{w/v}$) [50, 49, 43]. The experiments on BLGA and poly-L-lysine were performed at pH=7.0=7.0, with BLGA at a concentration $13~\mathrm{g\,L^{-1}}$ in the presence of $50~\mathrm{mM}$ NaCl, and $50~\mathrm{kDa}$ poly-L-lysine at concentration $5.4~\mathrm{g\,L^{-1}}$ with $100~\mathrm{mM}$ NaCl.
@@
Lysozyme $0.03090$ $1.30$ $0.25$
BLGA $0.01289$ $1.379$ 1.1-1.1
Poly-L-lysine $0.01296$ $1.324$ 1.1-1.1
Table 1: Parameters for the temperature dependence of the Soret coefficient $S_{T}^{\prime}$ [43] [left] and $S_{T}$ from our model [right] for different polypeptide solutions.
@@
Figure 2: Soret coefficient comparison between experiment data, empirical form, and model predictions shown for Lysozyme, BLGA, and Poly-L-Lysine. The experimental data shown have been extracted from Fig.1 of the reference [43] using an online data-digitization tool (PlotDigitizer).
@@
To elucidate the major role of the parameter $c$ in determining the shape of the $S_{T}$ curve, we show in Fig.3(a) and (b) the dependence of $\eta(T)$ and the corresponding $D_{0}(T)$ on temperature for all three cases. We have also plotted $\eta_{w}(T)$ (water) in these figures for a visual comparison. In Fig.3(c), we plot the derivative of $\ln D_{0}(T)$ with respect to temperature for all cases, including water. Fig.3(c) makes it clear that the first term on the right-hand side of eq.(17) is negative, while the second term is positive because $\beta>1$. From these curves, it is clear that while BLGA and Poly-L-Lysine are similar systems, Lysozyme is quite different in nature. Moreover, while the phoretic activity of Lysozyme is moderately dependent on the thermal expansion of water ($c=0.25$), the BLGA and Poly-L-Lysine show a substantial contribution of the correction to dynamic viscosity depending on the local thermal expansion coefficient of water $\alpha_{w}(T)$ with a c=1.1c=-$1.$ Note that there is a deviation of our curves from that of the empirical $S^{\prime}_{T}$ in the high temperature region for the BLGA and Poly-L-lysine. While fixing the parameter $c$, we wanted to keep it at a value for which the thermophilic-thermophobic transitions happen at the same temperature as those of $S^{\prime}_{T}$. We have realised that, in the high-temperature region, matching our graphs with the $S^{\prime}_{T}$ of BLGA and Poly-L-Lysine requires corrections to $\eta(T)$ beyond the term $c\alpha_{\omega}(T)$. Most probably, the coupling between the solvation mechanism and the diffusion in this range of temperature is much more complex than that captured by $c\alpha_{w}(T)$. Therefore, higher-order correction terms need to be taken in this high-temperature regime.
@@
To obtain a clearer estimate of the competing Fickian, non-Fickian diffusion currents and the solvation force-driven drift, which are in balance in the steady state, we show plots of these currents for each case under consideration against temperature in Fig.4. In these figures, one can see that the BLGA and Poly-L-lysine have very similar current density profiles, whereas Lysozyme shows a very different profile. Moreover, the dominant role in the competition of currents is played between the non-Fickian current and the drift induced by the solvation forces in the case of Lysozyme. However, for BLGA and Poly-L-Lysine, the dominant competition is between the Fickian and non-Fickian currents. The non-Fickian diffusion current for Lysozyme follows the general trend of the solvation current, i.e., it increases in magnitude with increasing temperature. The Fickian diffusion current generally decreases with increasing temperature, then increases. This makes the Fickian current cross zero near about $22\degree\text{C}$ for BLGA and Poly-L-Lysine, and near $25\degree\text{C}$ for Lysozyme. This corresponds to a maximum of $C(T)$.
@@
From ref.[49, 51], we find that a temperature difference of $\Delta T\sim 1\,\mathrm{K}$ is imposed over a plate length scale $\Delta x=0.7\,\mathrm{mm}$, yielding a temperature gradient $\Delta T/\Delta x\approx 10^{3}\,\mathrm{K\,m^{-1}}$. We have taken this value for $dT/dx$ in our calculations of current shown in Fig. 4. In all these figures, the total current is zero, as shown by the dotted central line. Fig.5(a) and (b) show the plot of concentration $C(T)$ and $dC(T)/dT$ against $T$ to present a visual guide as to where the Fickian current crosses the zero. Note that we have plotted concentrations only within the temperature range of the experiment. The temperature at the distribution’s maximum corresponds to the transition temperature, which must happen. The transition temperature is the one at which the Soret coefficient vanishes. This can also be confirmed by comparing Fig. 5(b), where the variation of the temperature derivative of Concentration with temperature is shown.
@@
Figure 5: Figure shows variation of (a) concentration (normalised), (b) derivative vs temperature plot obtained from our model for thermophoresis of Lysozyme, BLGA, Poly-l-lysine, respectively. In (c), we display the Comparison of normalised saturation concentration (normalised) with temperature obtained from equilibrium experiment and the model, shown for chicken egg-white lysozyme in the main body figure[52, 53, 49, 54, 50]. The upper-right inset (a semi-log plot) shows the variation in saturated concentration (un-normalised) with temperature.
@@
In Refs.[52, 49, 55, 56], Pusey and coworkers measured the saturation concentration $C_{\mathrm{sat}}$ of lysozyme solutions in equilibrium with tetragonal lysozyme crystals, which form a stable ordered phase below $25\,^{\circ}\mathrm{C}$. Measurements were performed for various pH values, bath temperatures, and salt concentrations. The part of the experiment relevant to our discussion is carried out at a fixed pH and temperature at a fixed (low) salt concentration. In this protocol, two columns containing lysozyme crystals are maintained in contact with a thermal bath at a temperature $T$. A solution consisting of buffer, water, NaCl, and lysozyme is passed through each column from one end. In the first column, an undersaturated solution is used, leading to dissolution of the solid lysozyme until the concentration of the solution exiting the column reaches a steady saturation value $C_{\mathrm{sat}}$. In the second column, a supersaturated solution is introduced, causing lysozyme from the solution to deposit onto the crystal until the same equilibrium saturation concentration $C_{\mathrm{sat}}$ is reached. In both cases, the system relaxes to a unique concentration at which dissolved lysozyme and solid crystals coexist in equilibrium, and this value is recorded as $C_{\mathrm{sat}}$.
@@
The experiment is then repeated at a different bath temperature, while keeping the pH and buffer properties unchanged. In this way, the temperature dependence of the saturation concentration is obtained. Experimentally, this dependence is well described by the empirical form given by Piazza et al. as $C_{\mathrm{sat}}(T)=C_{0}\exp\!\left(\frac{T}{T_{1}}\right),$ with $T_{1}$ in the range $8-13\,{\mathrm{K}}$  [52, 53, 49, 54, 50]. Importantly, no temperature gradient is present in these experiments, so it probes an equilibrium distribution rather than a thermophoretic steady state. The normalised experimental data of Pusey et al. at pH of 4.6, 3% NaCl in solution, and that of empirical form are compared with the normalised equilibrium saturation distribution called $C^{*}_{sat}$ predicted by our model, as shown in  5(c). We have chosen this data because it is closest to the thermophoresis experiment on Lysozyme, which was performed at pH 4.65 and about 2.33% NaCl. On comparison, we find an excellent match between our model and both data sets. The value of $T_{1}$ found for the empirical form is $T_{1}=9.95\,\mathrm{K}$. The inset in the same figure shows a semi-logarithmic plot of the saturation concentration $C_{\mathrm{sat}}$ (in units of g/mL) from the experimental data, the empirical form, and the prediction of our model. Note that we have set $a=0$ and $c=0$ in the strong concentration regime ($\beta$ is the same as that of Lysozyme in Fig. 5(a)) for our graph in Fig. 5(c).
@@
Figure 6: Soret coefficient obtained from the model keeping $c=0.25$ fixed and varying $\beta$ and $a$ shown in $(a)$ and variation of the corresponding concentration and its temperature derivative with temperature is shown in $(b)$ and $(c)$ respectively for the temperature range of $4\,^{\circ}\mathrm{C}$ to $38\,^{\circ}\mathrm{C}$ .
In Fig.6, we show some plots of the Soret coefficient and respective concentration obtained from our model for low concentration and high concentration diffusion regimes. Our model clearly shows that there can exist regimes in which the system remains completely thermophobic or thermophilic, or undergoes a transition from one regime to another, depending on the solvation parameter $a$, the exponent $\beta$, or the parameter $c$. Note that when the diffusivity has no explicit concentration-dependence ($\beta=0$), the Soret coefficient becomes $S_{T}=-a+\frac{d}{dT}\log D_{0}(T)$. The first term ‘$a$’ is the solvation term whose strength will be determined by the interaction of the particle surface with the fluid. When the contribution from the solvation term is negligible in comparison to the diffusive term, the Soret coefficient remains positive throughout in accordance with our model, as shown by parameter $a=0.01$ in Fig. 6(a). This kind of behaviour was quantitatively seen by Ludwig and Soret in their experiment in [20, 21, 57]. However, when the diffusive and solvation terms become comparable at some temperature, there can be a transition from thermophobic to thermophilic nature as shown in Fig. 6(a) using in the plot with $a=0.0309$. This transition is further confirmed by the concentration–temperature profile in Fig. 6(b), where the transition temperature corresponds to a local minimum in the concentration. Consistently, in Fig. 6(c) the derivative of the concentration with respect to temperature crosses zero at the same temperature. Moreover, it is worth pointing out that there also exists a regime (strong concentration-dependent diffusivity, here for $\beta=1.3$) in which the Soret coefficient can remain negative/positive throughout, as shown in Fig. 6(a). One can move these curves vertically by changing the solvation parameter $a$ to get a transition. The concentration and its temperature derivative corresponding to the parameters are displayed in Fig. 6(b) and (c), respectively.

### v2

@@ Abstract
In 1928, Chapman generalised Einstein’s theory of diffusion for non-uniform fluids to show the presence of a non-Fickian diffusion current, which he considered important in thermodiffusion (Ludwig-Soret effect). In 1941, Kiyosi Itô proposed the formal methods of stochastic calculus in the presence of spatially dependent diffusion, yielding the same non-Fickian diffusion current as shown by Chapman. The phenomenon of thermodiffusion and thermophoresis happens in the presence of a temperature gradient, which makes diffusion space-dependent. The role of solvation forces in thermophoresis will only be clearer once that of diffusion is understood properly. In this paper, we investigate the importance of Chapman’s non-Fickian diffusion current on the thermophoretic motion of colloidal particles in water (with weak salt concentration). We show that all the general features of variations of the Soret coefficient $S_{T}$ with temperature can be captured using Chapman’s non-Fickian diffusion current. We compare our theoretical results with experimental plots of the Soret coefficients for three polypeptides in aqueous solution: Lysozyme, BLGA, and Poly-L-Lysine, and find a strong match. We emphasise that, in addition to the as-yet-understood details of solvation forces, Chapman’s non-Fickian diffusion current is an indispensable element that must be taken into account for a complete understanding of thermophoresis and thermodiffusion.
Thermophoresis, Coordinate-dependent damping, Soret coefficient, Colloidal proteins, Brownian Motion
@@ I INTRODUCTION
The thermophoretic motion of colloidal particles is closely related to thermodiffusion, first reported by Ludwig in 1859 [20] and subsequently observed by Soret in 1879 [21]. The process of thermodiffusion, also known as the Ludwig-Soret effect, involves partial segregation of components in a solution or fluid mixture in the presence of a thermal gradient. Thermal diffusion is understood to play a significant role in the convective separation of components in a fluid mixture. It is the interplay of convective flow and diffusion that underlie the thermodiffusive partial separation of the mixture’s components. Therefore, a proper theoretical understanding of these complex processes of thermodiffusion and thermophoresis could only have meaningfully begun not before the advent of the physical theory of diffusion based on Einstein’s work on Brownian motion in 1905 [22, 23]. As is well acknowledged, Einstein’s theory of Brownian motion provides the microscopic origin of Fickian diffusion. What lacks wider acknowledgement is that Einstein’s theory of Brownian motion also underlies the non-Fickian diffusion in inhomogeneous media and fields.
@@
The Fick’s law of diffusion in a homogeneous medium gives the Fickian diffusion current density (e.g., in one dimension along the $x$-direction), written as $j_{F}=-D(dC(x)/dx)$ where the diffusion constant is $D$. Diffusivity is a constant that indicates the space in which diffusion takes place is uniform. The non-Fickian diffusion current density $j_{NF}$ will arise in an inhomogeneous medium where the diffusivity $D(x)$ is a function of space. This current will arise as an additive part of the total diffusion current density j=dD(x)C(x)/dxj=-dD(x)C(x)/dx, where $j_{NF}=-C(x)(dD(x)/dx)$. At the origin of both the parts of the diffusion current $j_{F}$ and $j_{NF}$, the assumption of the local homogeneity and isotropy of space, which is a fundamental requirement of Einstein’s theory of Brownian motion, remains intact [28]. This is why these are diffusion currents, unlike drift currents that break the local isotropy of space. From various perspectives, many groups have studied enhanced diffusion in active and other systems like enzymes and also in equilibrium Brownian systems, where this Chapman-Itô non-Fickian diffusion current has been taken into account [29, 30, 31, 32, 33, 28, 34, 35, 36, 37, 38, 39, 40]. Works by Di Pu et al. have considered the Chapman-Itô non-Fickian contribution in the field of thermophoresis  [41, 42, 43, 44]. Seuwin et al. [45] have tried to explain the sign change in protein thermophoresis from a hydrogen-bonding perspective, to cite a few. Recent experimental studies have confirmed the presence of Non-Fickian diffusion within assemblies of proteins [46].
@@
Keeping this general picture in mind, the process of thermophoresis (and thermodiffusion) can be looked at as a phenomenon arising from the competition of three current densities. (1) The Fickian diffusion current, (2) the Non-Fickian diffusion current, and (3) the drift current due to solvation forces. In the present paper, we demonstrate the interplay of these three competing ingredients, resulting in Soret coefficients observed in three distinct experiments on aqueous solutions of Lysozyme, BLGA, and Poly-L-Lysine. We chose these three experimental results to avoid cases involving strong electrostatic forces, which we will address in the future. We show that non-Fickian diffusion is a major contributor to thermophoresis and thermodiffusion. For the class of systems under consideration, we show that the local density of water (assumed to have low salt concentration), its thermal expansion coefficient, and a generic local pressure gradient can account for the observed $S_{T}(T)$ of all the test cases. We also show that our model can match the experimentally observed equilibrium saturation concentration curve for lysozyme quite well, underscoring the important role the non-Fickian current plays in equilibrium and nonequilibrium phenomena. Moreover, we show that the theoretical model built with these specific processes of thermophoresis in colloids in mind can also account for the general features of the Soret coefficient for thermodiffusion. Therefore, the model framework developed in this paper, based on results from kinetic theory and empirical inputs, may capture some fundamental aspects of transport driven by thermal gradients in a fluid.
@@ II.1 The general structure of the diffusivity
The $D_{0}(T)$ represents the bare single-particle diffusivity, capturing the explicit temperature dependence arising from solute–solvent interactions (e.g. viscosity, hydration, and thermal agitation) in the dilute limit. The concentration-dependent factor of diffusivity $\Lambda(C)$ accounts for the dependence of diffusivity on crowding and ensuing hydrodynamic effects in general [47, 48, 49, 50]. The parameter $b$ sets the strength of the concentration dependence of the diffusivity. One may imagine that there exists a scale of concentration $C_{0}$, where $b=C_{0}^{-\beta}$. The dependence of the diffusivity on the local concentration is dominant for $C>C_{0}$.
@@
At this stage, we can define the dominant asymptotic regimes of the concentration dependence of the diffusivity as (1) the weak concentration $bC^{\beta}<<1$ and (2) the strong concentration $bC^{\beta}>>1$, where the corresponding Soret coefficients at the leading order are, respectively,
@@ II.2 Temperature dependence of the dynamic viscosity
Thermal gradients can alter the local solvent density, thereby modifying the local dynamic viscosity through a range of processes, e.g., from local hydrodynamics to hydrogen-bonding chemistry. The strength of this viscosity modification due to temperature gradient can be quantified by the solvent thermal expansion coefficient $\alpha_{w}(T)$, defined in terms of the temperature-dependent solvent density $\rho_{w}(T)$ as
@@ III Results: Comparison with experiment
Let us compare our model with the curves fitted to the experimental data for hen egg-white lysozyme and other polypeptide solutions reported in ref. [51, 57, 58]. The empirical fitting function chosen to fit the experimentally obtained data for the Soret coefficient is
@@
where $S^{\infty}_{T}$ is the high-temperature thermophobic limit, $T^{*}$ is the temperature at which $S^{\prime}_{T}$ switches sign, and $T_{0}$ denotes the scale of exponential growth. Note that the transition temperature $T^{*}$ is poorly sensitive to salt addition and Lysozyme charge [51, 57, 58]. The above-mentioned empirical function for the Soret coefficient in these systems is applicable over a wide range of pH and ionic strength, indicating that neither pH nor ionic strength significantly affects the ansatz. Our model at present does not include explicit contributions from pH or ionic strength; beyond this limit, they are accounted for through the parameters $c$ and $a$. Thus, our model is well-suited for comparison with experimental data that, in general, matches the empirical Soret coefficient of eq.(16). The value of the three free parameters of $S_{T}^{\prime}$ in eq.(16) and the corresponding three free parameters in the expression of $S_{T}$ in our model, which make the curves match $S^{\prime}_{T}$ that matches the experimental data, are shown in Table 1. Figure 2 compares curves with the $S_{T}$ from our model plotted as red dots. The experiments on lysozyme, whose data are presented here, were carried out in a sodium acetate buffer at pH=4.65=4.65, with a lysozyme concentration $7~\mathrm{g\,L^{-1}}$ and $0.4~\mathrm{M}$ NaCl (corresponding to $2.3\%~\mathrm{w/v}$) [58, 57, 51]. The experiments on BLGA and Poly-L-lysine were performed at pH=7.0=7.0, with BLGA at a concentration $13~\mathrm{g\,L^{-1}}$ in the presence of $50~\mathrm{mM}$ NaCl, and $50~\mathrm{kDa}$ Poly-L-Lysine at concentration $5.4~\mathrm{g\,L^{-1}}$ with $100~\mathrm{mM}$ NaCl.
@@
Lysozyme $0.0314$ $1.265$ $0.275$
BLGA $0.0303$ $1.232$ $0.183$
Poly-L-Lysine $0.0303$ $1.179$ $0.195$
Table 1: Parameters for the temperature dependence of the Soret coefficient $S_{T}^{\prime}$ [51] [left] and $S_{T}$ from our model [right] for different polypeptide solutions obtained via least-squares minimisation with respect to experimental data; the details can be found in the Appendix.
@@
Figure 2: Soret coefficient comparison between experiment data, empirical form, and model predictions shown for Lysozyme, BLGA, and Poly-L-Lysine. The experimental data shown have been extracted from Fig.1 of the reference [51] using an online data-digitisation tool (PlotDigitizer).
@@
To elucidate the major role of the parameter $c$ in determining the shape of the $S_{T}$ curve, we show in Fig. 3(a) and  3(b) the dependence of $\eta(T)$ and the corresponding $D_{0}(T)$ on temperature for all three cases. We have also plotted $\eta_{w}(T)$ (water) in these figures for a visual comparison. In Fig. 3(c), we plot the derivative of $\ln D_{0}(T)$ with respect to temperature for all cases, including water. Fig. 3(c) makes it clear that the first term on the right-hand side of eq.(17) is negative, while the second term is positive because $\beta>1$. From these curves, it is clear that while BLGA and Poly-L-Lysine are similar systems, Lysozyme is somewhat a different system in nature. Moreover, while the phoretic activity of Lysozyme is substantially dependent on the thermal expansion of water ($c=0.275$), the BLGA and Poly-L-Lysine show a relatively reduced contribution of the correction to dynamic viscosity depending on the local thermal expansion coefficient of water $\alpha_{w}(T)$ with a $c\simeq 0.183$ and $c\simeq 0.195$, respectively. Since $\alpha_{w}$ is practically positive in the considered temperature range, the coupling between the solvation mechanism and diffusion in this temperature range, due to a positive $c$, effectively reduces viscosity in the solvation layer. The physics of this, which the present model indicates, hides in the actual solvation mechanism at play, which needs to be understood.
@@
To obtain a clearer estimate of the competing Fickian, non-Fickian diffusion currents and the solvation force-driven drift, which are in balance in the steady state, we show plots of these currents for each case under consideration against temperature in Fig. 4. In these figures, one can see that the BLGA and Poly-L-lysine have very similar current density profiles, whereas Lysozyme shows a distinct profile, with the rise in the solvation current with temperature occurring more at the higher temperature end. Moreover, the dominant role in the competition of currents is played between the non-Fickian current and the drift induced by the solvation forces. The Fickian diffusion current is smaller in magnitude than the other two currents; however, it reverses direction when the corresponding concentration gradient changes sign with temperature, thereby changing the concentration gradient in space in the presence of a linear temperature gradient. This makes the Fickian current cross zero near about $22\degree\text{C}$ for BLGA and Poly-L-Lysine, and near $25\degree\text{C}$ for Lysozyme, as shown in Fig. 5(a).
@@
From ref. [57, 59], we find that a temperature difference of $\Delta T\sim 1\,\mathrm{K}$ is imposed over a plate length scale $\Delta x=0.7\,\mathrm{mm}$, yielding a temperature gradient $\Delta T/\Delta x\approx 10^{3}\,\mathrm{K\,m^{-1}}$. We have taken this value for $dT/dx$ in our calculations of current shown in Fig. 4. In all these figures, the total current is zero, as shown by the dotted central line. Fig. 5(a) and Fig. 5(b) show the plot of concentration $C(T)$ and $dC(T)/dT$ against $T$ to present a visual guide as to where the Fickian current crosses the zero. Note that we have plotted concentrations only within the temperature range of the experiment. The temperature at the distribution’s maximum corresponds to the transition temperature, which must happen. The transition temperature is the one at which the Soret coefficient vanishes. This can also be confirmed by comparing Fig. 5(b), where the variation of the temperature derivative of Concentration with temperature is shown.
@@
Figure 5: Figure shows variation of (a) concentration (normalised), (b) derivative vs temperature plot obtained from our model for thermophoresis of Lysozyme, BLGA, Poly-L-Lysine, respectively. In (c), we display the Comparison of normalised saturation concentration ($C^{*}_{sat}$) with temperature obtained from equilibrium experiment and the model, shown for chicken egg-white lysozyme in the main body figure [60, 61, 57, 62, 58]. The upper-right inset (a semi-log plot) shows the variation in saturated concentration (un-normalised) with temperature.
@@
Concentration-dependent cluster formation in lysozyme solutions and non-Fickian diffusion in proteins have been observed in experiments [63, 46]. In Refs.[60, 57, 64, 65], Pusey and coworkers measured the saturation concentration $C_{\mathrm{sat}}$ of lysozyme solutions in equilibrium with tetragonal lysozyme crystals, which form a stable ordered phase below $25\,^{\circ}\mathrm{C}$. Measurements were performed for various pH values, bath temperatures, and salt concentrations. The part of the experiment relevant to our discussion is carried out at a fixed pH and temperature at a fixed (low) salt concentration. In this protocol, two columns containing lysozyme crystals are maintained in contact with a thermal bath at a temperature $T$. A solution consisting of buffer, water, NaCl, and lysozyme is passed through each column from one end. In the first column, an undersaturated solution is used, leading to dissolution of the solid lysozyme until the concentration of the solution exiting the column reaches a steady saturation value $C_{\mathrm{sat}}$. In the second column, a supersaturated solution is introduced, causing lysozyme from the solution to deposit onto the crystal until the same equilibrium saturation concentration $C_{\mathrm{sat}}$ is reached. In both cases, the system relaxes to a unique concentration at which dissolved lysozyme and solid crystals coexist in equilibrium, and this value is recorded as $C_{\mathrm{sat}}$.
@@
The experiment is then repeated at a different bath temperature, while keeping the pH and buffer properties unchanged. In this way, the temperature dependence of the saturation concentration is obtained. Experimentally, this dependence is well described by the empirical form given by Piazza et al. as $C_{\mathrm{sat}}(T)=C_{0}\exp\!\left(\frac{T}{T_{1}}\right),$ with $T_{1}$ in the range $8-13\,{\mathrm{K}}$  [60, 61, 57, 62, 58]. Importantly, no temperature gradient is present in these experiments, so it probes an equilibrium distribution rather than a thermophoretic steady state. The normalised experimental data of Pusey et al. at pH of 4.6, 3% NaCl in solution, and that of empirical form are compared with the normalised equilibrium saturation distribution called $C^{*}_{sat}$ predicted by our model, as shown in Fig. 5(c). We have chosen this data because it is closest to the thermophoresis experiment on Lysozyme, which was performed at pH 4.65 and about 2.33% NaCl. On comparison, we find an excellent match between our model and both data sets. The value of $T_{1}$ found for the empirical form is $T_{1}=9.95\,\mathrm{K}$. The inset in the same figure shows a semi-logarithmic plot of the saturation concentration $C_{\mathrm{sat}}$ (in units of g/mL) from the experimental data, the empirical form, and the prediction of our model. Note that we have set $a=0$ and $c=0$ in the strong concentration regime ($\beta$ is the same as that of Lysozyme in Fig. 5(a)) for our graph in Fig. 5(c).
@@
Figure 6: Soret coefficient obtained from the model keeping $c=0.275$ fixed and varying $\beta$ and $a$ shown in $(a)$ and variation of the corresponding concentration and its temperature derivative with temperature is shown in $(b)$ and $(c)$ respectively for the temperature range of $4\,^{\circ}\mathrm{C}$ to $38\,^{\circ}\mathrm{C}$ .
In Fig. 6, we show some plots of the Soret coefficient and respective concentration obtained from our model for low concentration and high concentration diffusion regimes. Our model clearly shows that there can exist regimes in which the system remains completely thermophobic or thermophilic, or undergoes a transition from one regime to another, depending on the solvation parameter $a$, the exponent $\beta$, or the parameter $c$. Note that when the diffusivity has no explicit concentration-dependence ($\beta=0$), the Soret coefficient becomes $S_{T}=-a+\frac{d}{dT}\log D_{0}(T)$. The first term ‘$a$’ is the solvation term whose strength will be determined by the interaction of the particle surface with the fluid. When the contribution from the solvation term is negligible in comparison to the diffusive term, the Soret coefficient remains positive throughout in accordance with our model, as shown by parameter $a=0.01$ in Fig. 6(a). This kind of behaviour was quantitatively seen by Ludwig and Soret in their experiment in [20, 21, 66]. However, when the diffusive and solvation terms become comparable at some temperature, there can be a transition from thermophobic to thermophilic nature as shown in Fig. 6(a) using in the plot with $a=0.0303$. This transition is further confirmed by the concentration–temperature profile in Fig. 6(b), where the transition temperature corresponds to a local minimum in the concentration. Consistently, in Fig. 6(c) the derivative of the concentration with respect to temperature crosses zero at the same temperature. Moreover, it is worth pointing out that there also exists a regime (strong concentration-dependent diffusivity, here for $\beta=1.23$) in which the Soret coefficient can remain negative/positive throughout, as shown in Fig. 6(a). One can move these curves vertically by changing the solvation parameter $a$ to get a transition. The concentration and its temperature derivative corresponding to the parameters are displayed in Fig. 6(b) and Fig. 6(c), respectively.